# Distributed acoustic sensing for ocean applications

### v2

@@ II.2 Rayleigh backscattered wave
The total backscattered field at any time instant $\tau$ results from the superposition of the backscattered light from all scatterers within the isochronous illuminated fiber length defined by the pulse length $L_{p}=T_{p}c_{n}$, which making use of (6), (2), and (3) yields
$E_{b}(\tau)$ $\!=\!\!\int\limits_{-\infty}^{\infty}\!\!E_{s}(\tau|z_{s})\mathrm{d}z_{s}=\!\!\int\limits_{-\infty}^{\infty}\!\!E_{p}\left(\tau\!-\!2\frac{z_{s}}{c_{n}}\right)s(z_{s})\mathrm{d}z_{s}$ (8)
$\!=\!\!\!\!\int\limits_{-\infty}^{\infty}\!\!p\!\left(\!\frac{\tau\!-\!2\frac{z_{s}}{c_{n}}}{T_{p}}\!\right)\!\cos\!\left(\!\omega\!\left(\!\tau\!-\!2\frac{z_{s}}{c_{n}}\right)\!\!+\!\phi_{0}\!\right)\!\!s(z_{s})\mathrm{d}z_{s}.$
The received time signal of the total backscattered field is mapped to the fiber’s axial dimension as [Fig. 2(h)]:
$E_{b}(z)=\!\!\!\!\int\limits_{-L_{p}/2}^{L_{p}/2}\!\!\!\!w(z-z_{s})\cos\left(2k(z\!-\!z_{s})\!+\!\phi_{0}\right)s(z_{s})\mathrm{d}z_{s},$ (9)
where $k=\omega/c_{n}$ is the wavenumber and $p\left((\tau-t_{s})/T_{p}\right)=w(\tau-t_{s})$ for $\lvert\tau-t_{s}\rvert\leq T_{p}/2$ is equivalent to $p\left((z-z_{s})/L_{p}\right)=w(z-z_{s})$ for $\lvert z-z_{s}\rvert\leq L_{p}/2$ with a linear transformation of variables between time and space through the speed of light, $\tau=2z/c_{n}$. The resulting backscattered field (9) is a stochastic process resulting from the convolution of the pulse smoothing kernel with the scattering strength $s$, which is a random function of the axial dimension $z$.
@@ II.3 Spectral analysis of the backscattered field
and the spatial Fourier transform is obtained by replacing $\omega=kc_{n}$ and $\tau=2z/c_{n}$
@@
Then, the probe pulse (2) is expressed as
$\widetilde{E}_{p}(t,\omega)=p\left(\frac{t}{T_{p}}\right)e^{j(\omega t+\phi_{0})}.$ (12)
Consequently, the backscattered wave from a point scatterer (6) is
$\widetilde{E}_{s}(\tau,\omega|z_{s})$ $=\widetilde{E}_{p}\left(\tau-2\frac{z_{s}}{c_{n}},\omega\right)s(z_{s})$ (13)
$=p\left(\frac{\tau-2\frac{z_{s}}{c_{n}}}{T_{p}}\right)e^{j(\omega(\tau-2\frac{z_{s}}{c_{n}})+\phi_{0})}s(z_{s})$
and the total backscattered field due to the spatial distribution of scatterers within a pulse length (8) is
$\widetilde{E}_{b}(\tau,\omega)$ $=\int\limits_{-\infty}^{\infty}\widetilde{E}_{s}(\tau,\omega|z_{s})\mathrm{d}z_{s}$ (14)
$=\!\!\int\limits_{-\infty}^{\infty}\!\!p\bigg{(}\frac{\tau\!-\!2\frac{z_{s}}{c_{n}}}{T_{p}}\bigg{)}\!e^{j\left(\omega\left(\tau\!-\!2\frac{z_{s}}{c_{n}}\right)+\phi_{0}\right)}s(z_{s})\mathrm{d}z_{s}.$
It follows that the total backscattered field expressed in the spatial dimension (9) is,
$\widetilde{E}_{b}(z,k)$ $=e^{j(2kz+\phi_{0})}\!\!\!\!\int\limits_{-L_{p}/2}^{L_{p}/2}\!\!\!\!w(z-z_{s})s(z_{s})e^{-j2kz_{s}}\mathrm{d}z_{s}$ (15)
$=e^{j(2kz+\phi_{0})}\left[\widetilde{W}(k)\widetilde{S}(k)\right],$
where $\widetilde{W}(k)$ and $\widetilde{S}(k)$ denote the spatial Fourier transform of the pulse window and the scattering function, respectively. For example, considering the pulse window to be a rectangular function, $\widetilde{W}(k)$ is a sinc function as detailed later in  (24). The scattering function $s(z_{s})$ is a random process with zero mean and autocorrelation function $R_{ss}$. Assuming that the scatterers in the fiber are spatially uncorrelated, $R_{ss}(z_{s})=\sigma^{2}\delta\!\left(z_{s}\right)$. Hence, the Fourier transform of the scattering function is a random process with constant spectral density as results from the Fourier transform of the autocorrelation function, $\widetilde{C}_{s}(k)=\mathcal{F}(R_{ss}(z_{s}))=\sigma^{2}$. The time-frequency (14) or, equivalently, the space-wavenumber (15) formulation expresses the total backscattered field as the short-time (spatial) Fourier transformAllen1977 of the scattering function over a window defined by the probe pulse.
@@ II.4 Differential phase-measuring coherent optical reflectometry
DAS systems employ phase-sensitive coherent optical reflectometry methods (see Appendix) to measure the phase shift in the backscattered light from two fiber sections [Fig. 2(h)]. These differential phase measurements are a proxy for estimating strain from mechanical vibrations [Hartog2017, , Ch. 6]. There are several implementations of differential phase-measuring methods. Still, the underlying principle is mixing the backscattered signal with itself at a relative delay and detecting the phase of the resulting cross-correlation. The two signals can be mixed either in the optical domain (direct detection techniques) or in the electrical domain (coherent detection techniques) and can have the same frequency (interferometric recovery technique and homodyne detections) or slightly different frequencies (dual-pulse systems or heterodyne detection).Hartog2017 ; He2021
In any case, the resulting interferometric field is detected with coherent optical reflectometry methods (36) by mixing the backscattered light $\widetilde{E}_{b}(\tau,\omega)$ with the corresponding backscattered light $\widetilde{E}_{b}(\tau+\Delta\tau,\omega)$ at a relative delay of $\Delta\tau$. For an undisturbed fiber, the relative delay corresponds to a fixed distance along the fiber referred to as the gauge length $L_{g}$. External mechanical vibrations deform the optical fiber cable and, consequently, perturb the gauge length by $\Delta L$ resulting in an actual relative delay of $\Delta\tau=2(L_{g}+\Delta L)/c_{n}$. Hence, the interference product at the photodetector is [Hartog2017, , Eq. (6.11)]
$\widetilde{E}_{b}\left(\tau,\omega\right)\widetilde{E}^{*}_{b}\left(\tau+\Delta\tau,\omega\right)=$ (16)
$\quad\Big{(}\int\limits_{-\infty}^{\infty}\!\!p\Big{(}\frac{\tau\!-\!2\frac{z_{s}}{c_{n}}}{T_{p}}\Big{)}\!e^{j\left(\omega\left(\tau\!-\!2\frac{z_{s}}{c_{n}}\right)+\phi_{0}\right)}s(z_{s})\mathrm{d}z_{s}\Big{)}\cdot$
$\Big{(}\!\!\int\limits_{-\infty}^{\infty}\!\!p\Big{(}\frac{\tau\!+\!2\frac{L_{g}+\Delta L-z_{s}}{c_{n}}}{T_{p}}\Big{)}\!e^{-j\left(\omega\left(\tau\!+\!2\frac{L_{g}+\Delta L-z_{s}}{c_{n}}\right)+\phi_{0}\right)}s(z_{s})\mathrm{d}z_{s}\!\!\Big{)},$
or mapped in the space-wavenumber domain,
$\widetilde{E}_{b}\left(z,k\right)\widetilde{E}^{*}_{b}\left(z+L_{g}+\Delta L,k\right)=$ (17)
$\quad\Big{(}\!e^{j(2kz+\phi_{0})}\!\!\!\!\!\int\limits_{-L_{p}/2}^{L_{p}/2}\!\!\!\!w(z\!-\!z_{s})s(z_{s})e^{-j2kz_{s}}\mathrm{d}z_{s}\!\Big{)}\cdot$
$\quad\Big{(}\!e^{-j(2k(z+L_{g}+\Delta L)+\phi_{0})}\!\!\!\!\!\int\limits_{-L_{p}/2}^{L_{p}/2}\!\!\!\!w(z\!-\!z_{s})s(z_{s})e^{j2kz_{s}}\mathrm{d}z_{s}\!\Big{)}$
$=e^{-jk(2\Delta L)}\left[\widetilde{W}(k)\widetilde{S}(k)\right]\left[\widetilde{W}(k)\widetilde{S}(k)e^{j2kL_{g}}\right]^{*}$
@@
where $\Delta\Phi=k(2\Delta L)$ and $\widetilde{C}_{b(L_{g})}(k)=\left[\widetilde{W}(k)\widetilde{S}(k)\right]\left[\widetilde{W}(k)\widetilde{S}(k)e^{j2kL_{g}}\right]^{*}$ denotes the cross-spectral density of the backscattered signal at a relative delay determined by the gauge length $L_{g}$.
The gauge length determines the amount of overlap of the backscattered signals in (16), i.e. $L_{\text{overlap}}=\max\left(L_{p}-L_{g},0\right)$. Hence, the mixed signals are correlated for a gauge length shorter than the pulse length, $L_{g}<L_{p}$, due to the autocorrelation of the pulse apodization function. The gauge length should be longer than the pulse length to measure the effect of phase modulation due to external vibrations independently of the pulse that probes the fiber.
Figure 3: Three-dimensional Cartesian coordinate system indicating the direction $\mathbf{e}$ of an impinging wavefield in terms of the azimuth angle $\phi$ and the elevation angle $\theta$. By convention, the linear optical fiber cable is aligned with the z-axis. The directional angles are detailed in the corresponding two-dimensional perspectives.
Figure 4: Two-dimensional data structure in DAS. (a) An interrogator uses a laser source to transmit a pulse train $E_{p}(t)$ into the optical fiber, and (b) receives the backscattered signal $E_{b}(t)$ due to inhomogeneities in the dielectric material. (c) Back at the laser’s position, the received time series is converted to slow time $t$ and fast time $\tau$. The time scales are determined by the acquisition sampling frequency $f_{s}$ along the fast-time axis $\tau$, and the pulse repetition frequency $f_{r}$ along the slow-time axis $t$. (d) Measurements along the fast-time correspond to the spatial dimension along the fiber, $z=\tau c_{n}/2$, and are commonly downsampled to the channel spacing $L_{h}$. Phase variations in the backscattered signal measured along distance $z$ are associated with the applied axial strain $\epsilon_{zz}$, whereas phase variations in the backscattered signal measured across time $t$ are associated with the strain rate $\partial\epsilon_{zz}/\partial t$.
@@ II.5 Axial strain estimates
The analysis herein shows that the differential phase measurements of the light backscatter are linearly related to the axial strain of the fiber, which, in turn, is related to acoustic quantities in the environment. The subscript $a$ is introduced to indicate acoustic quantities and differentiate them from the corresponding electromagnetic quantities in the previous analysis. Namely, $\omega_{a}$ denotes the radial frequency of the acoustic wave, $k_{a}=\omega_{a}/c_{a}$ denotes the wavenumber and $c_{a}$ denotes the sound speed.
Interrogating the optical fiber cable with phase-sensitive coherent optical reflectometry methods (17) provides a distributed measurement of the differential phase $\Delta\Phi$ caused by optical path length variations $\Delta L$. Specifically,
$\Delta\Phi$ $=2k\Delta L=2\frac{2\pi(n+\Delta n)}{\lambda}L_{g}\frac{\Delta L}{L_{g}}$ (18)
$=\frac{4\pi nL_{g}}{\lambda}(1+\frac{\Delta n}{n})\frac{\Delta L}{L_{g}}=\frac{4\pi n\xi L_{g}}{\lambda}\frac{\Delta L}{L_{g}},$
where $\lambda$ is the wavelength of the probe pulse and $\xi=1+\Delta n/n$ is the Pockels coefficient that accounts for the strain-optical effect, i.e. strain applied to the fiber both changes its length and modifies its refractive index.Hartog2017 ; williams2024toward In single-mode fiberglass the Pockels coefficient is $\xi=0.79$.Lindsey2021
Considering an optical fiber cable that is coupled with an elastic medium, e.g., deployed on the seafloor, the cable’s deformation $\Delta L$ is attributed to mechanical waves impinging on it [Fig. 2(e)]. Specifically, let $\widetilde{\mathbf{u}}(t,\mathbf{r})=(A\mathbf{e})\odot e^{j(\omega_{a}t-k_{a}\mathbf{e}\odot\mathbf{r})}$ denote the particle displacement vector, defined in three-dimensional space $\mathbf{r}=[x,y,z]^{T}$, due to an acoustic plane wave with amplitude $A$. The direction of the plane wave is defined by the azimuthal angle $\phi$ and the polar angle $\theta$ measured from the normal to the cable axial direction $z$ through the directional vector $\mathbf{e}=[\cos\phi\cos\theta,\sin\phi\cos\theta,\sin\theta]^{T}$ as depicted in Fig. 3. The $\odot$ operator denotes element-wise multiplication.
The particle displacement component along the $z$ direction, $\widetilde{u}_{z}(t,z)=A\sin\theta e^{j(\omega_{a}t-k_{a}\sin\theta z)}$, is related to the axial strain as [Ewing1957, , Eq. (1-2)]
@@
$\quad\quad\quad=\widetilde{u}_{z}(t,z)\frac{e^{-jk_{a}\sin\theta\frac{L_{g}}{2}}-e^{jk_{a}\sin\theta\frac{L_{g}}{2}}}{L_{g}}$
$\quad\quad\quad=\widetilde{u}_{z}(t,z)\frac{-j2\sin(k_{a}\sin\theta\frac{L_{g}}{2})}{L_{g}}$
$\quad\quad\quad=\widetilde{u}_{z}(t,z)\left(-jk_{a}\sin\theta\right)\frac{\sin(k_{a}\sin\theta\frac{L_{g}}{2})}{k_{a}\sin\theta\frac{L_{g}}{2}}$
$\quad\quad\quad=\!\!-jk_{a}\sin^{2}\!\theta Ae^{j(\omega_{a}t-k_{a}\!\sin\!\theta z)}\mathrm{sinc}(L_{g}k_{z}),$
where $k_{z}=k_{a}\sin\theta/(2\pi)$ and the spatial derivative is approximated by the total deformation over a finite gauge length.
Equation (18) indicates that the measured differential phase is linearly related to an estimate of the axial strain over the gauge length $\widehat{\epsilon}_{zz}=\Delta L/L_{g}$ as
$\widehat{\epsilon}_{zz}=\frac{\lambda}{4\pi n\xi L_{g}}\Delta\Phi.$ (20)
In turn, the axial strain estimate (20) is related to the particle displacement of external vibrations such as the acoustic wavefield around the optical-fiber cable through (19), which reveals two important aspects for interpreting DAS measurements.
First, the dependence of the axial strain amplitude to the square of the sine of the wavefield’s angle of incidence makes the strain measurement insensitive to broadside incidence, i.e., $\widetilde{\epsilon}_{zz}=0$ for $\theta=0$. This angular dependency is crucial in interpreting DAS data, as it influences the amplitude and detectability of acoustic signals. It also limits the effectiveness of DAS in configurations where ray paths are predominantly orthogonal to the fiber-optic cable. To overcome this limitation, helically wound cables were developed, offering broadside sensitivity and enabling the observation of signals from a wider range of angles.kuvshinov2016interaction ; hornman2017field
Second, the finite difference approximation of the spatial derivative of the displacement over the gauge length introduces directivity to the measurement defined by the term $\mathrm{sinc}(L_{g}k_{z})$. Hence, the gauge length acts as a spatial window which determines the measurement resolution; see Sec. III for details.
Some DAS systems, instead of estimating the axial strain (20) by unwrapping the differential phase measurement $\Delta\Phi$, estimate the strain rate by differentiating the phase measurement between successive pulses [Hartog2017, , Sec. 6.3.2], $\frac{\partial\epsilon_{zz}}{\partial t}$. In such systems, denoting the pulse repetition period as $T$, the axial strain rate estimate is
@@
$\approx(-jk_{a}\sin\theta)\mathrm{sinc}(L_{g}k_{z})\frac{\partial\widetilde{u}_{z}}{\partial t}$
$\approx\omega_{a}k_{a}\sin^{2}\theta Ae^{j(\omega_{a}t-k_{a}\sin\theta z)}\mathrm{sinc}(L_{g}k_{z}).$
Concisely, axial strain estimates are related to the particle displacement component along the fiber, whereas axial strain rate estimates are related to the corresponding component of particle velocity.

### v3

@@ II.2 Rayleigh backscattered wave
The total backscattered field received at $\tau=2z/c_{n}$ results from the superposition of the backscattered light from all scatterers within the illuminated fiber length at $t=z/c_{n}$ defined by the pulse length $L_{p}=T_{p}c_{n}$, since $p\left((t-t_{s})/T_{p}\right)=w(t-t_{s})$ for $\lvert t-t_{s}\rvert\leq T_{p}/2$ is equivalent to $p\left((z-z_{s})/L_{p}\right)=w(z-z_{s})$ for $\lvert z-z_{s}\rvert\leq L_{p}/2$ with a linear transformation of variables between time and space through the speed of light. Making use of Eqs. (9), (2),(3) and (5) yields
$E_{b}(\tau)$ $\!=\!\!\!\!\!\!\int\limits_{z-L_{p}/2}^{z+L_{p}/2}\!\!\!\!\!\!E_{s}(\tau|z_{s})\mathrm{d}z_{s}\!\!\!\ =\!\!\!\!\!\!\!\!\int\limits_{z-L_{p}/2}^{z+L_{p}/2}\!\!\!\!\!\!E_{p}\left(\tau\!-\!2\frac{z_{s}}{c_{n}}\right)s(z_{s})\mathrm{d}z_{s}$ (11)
$\!=\!\!\!\!\!\!\int\limits_{z-L_{p}/2}^{z+L_{p}/2}\!\!\!\!\!w\!\left(\!z\!-\!z_{s}\!\right)\!\cos\!\left(\!\omega_{p}\!\left(\!\tau\!-\!2\frac{z_{s}}{c_{n}}\right)\!\!+\!\phi_{0}\!\right)\!\!s(z_{s})\mathrm{d}z_{s}.$
Then, with wavenumber $k_{p}=\omega_{p}/c_{n}$, the received time signal of the total backscattered field is mapped to the spatial dimension along the fiber’s axis as [Fig. 2(h)]:
$E_{b}(z)$ $\!=\!\!\!\!\!\!\int\limits_{z-L_{p}/2}^{z+L_{p}/2}\!\!\!\!\!\!w(z-z_{s})\cos\left(2k_{p}(z\!-\!z_{s})\!+\!\phi_{0}\right)s(z_{s})\mathrm{d}z_{s}$ (12)
$=\left(E_{p}*s\right)(z).$
The resulting backscattered field in Eq. (12) is a stochastic process resulting from the convolution of the pulse smoothing kernel with the scattering strength $s$, which is a random function of the spatial dimension along the fiber’s axis $z$.
@@ II.3 Spectral analysis of the backscattered field
and the spatial Fourier transform is obtained by replacing $\omega=kc_{n}$ and $\tau=2z/c_{n}$,
@@
In complex notation, the probe pulse in Eq. (2) for monochromatic transmission in Eq. (5) is expressed as
$\widetilde{E}_{p}(t)=p\left(\frac{t}{T_{p}}\right)e^{j(\omega_{p}t+\phi_{0})},$ (15)
and in the frequency domain,
$\widetilde{E}_{p}(\omega)$ $=\int\limits_{-\infty}^{\infty}p\left(\frac{t}{T_{p}}\right)e^{j(\omega_{p}t+\phi_{0})}e^{-j\omega t}\mathrm{d}{t}$ (16)
$=\int\limits_{-T_{p}/2}^{T_{p}/2}w(t)e^{-j(\omega-\omega_{p})t}e^{j\phi_{0}}\mathrm{d}{t}$
$=\widetilde{W}(\omega-\omega_{p})e^{j\phi_{0}},$
where $\widetilde{W}(\omega-\omega_{p})$ denotes the temporal Fourier transform $\widetilde{W}(\omega)$ of the window function $w(t)$ centered at the angular frequency of the pulse $\omega_{p}$.
Similarly, the total backscattered field in Eq. (11) is expressed as
$\widetilde{E}_{b}(\tau)=e^{j\phi_{0}}\!\!\!\!\int\limits_{z-L_{p}/2}^{z+L_{p}/2}\!\!\!\!\!w\!\left(z\!-\!z_{s}\right)\!e^{j\omega_{p}\left(\tau\!-\!2\frac{z_{s}}{c_{n}}\right)}s(z_{s})\mathrm{d}z_{s},$ (17)
and in space from Eq. (12),
$\widetilde{E}_{b}(z)=e^{j\phi_{0}}\!\!\!\!\int\limits_{z-L_{p}/2}^{z+L_{p}/2}\!\!\!\!w(z-z_{s})e^{j2k_{p}(z-z_{s})}s(z_{s})\mathrm{d}z_{s}.$ (18)
With $\widetilde{W}(k)$ and $\widetilde{S}(k)$ denoting the spatial Fourier transform of the pulse window $w(z)$ and the scattering function $s(z)$, respectively, the spatial Fourier transform of the total backscattered field in Eq. (18) is,
$\widetilde{E}_{b}(k)\!\!=\!\!e^{j\phi_{0}}\!\!\left[\widetilde{W}(k\!-\!k_{p})\widetilde{S}(k)\right]\!\!=\!\!e^{j\phi_{0}}\!\!\left[\widetilde{W}(k)\widetilde{S}(k\!-\!k_{p})\right]\!.$ (19)
For example, considering the pulse window to be a rectangular function, $\widetilde{W}(k)$ is a sinc function, as detailed later in Eq. (30). The scattering function $s(z)$ is a random process with zero mean and autocorrelation function $R_{ss}$. Assuming that the scatterers in the fiber are spatially uncorrelated with variance $\sigma^{2}$, then $R_{ss}(z)=\sigma^{2}\delta\!\left(z\right)$. Hence, the Fourier transform of the scattering function is a random process with constant spectral density as results from the Fourier transform of the autocorrelation function $\widetilde{C}_{s}(k)=\mathcal{F}(R_{ss}(z))=\sigma^{2}$.
@@ II.4 Differential phase-measuring coherent optical reflectometry
DAS systems employ phase-sensitive coherent optical reflectometry methods (see Appendix) to measure the phase shift in the backscattered light from two fiber locations [Fig. 2(h)]. These differential phase measurements [Fig. 2(i)] are linearly related to strain from mechanical vibrations and are used as proxy to strain estimates [Hartog2017, , Ch. 6].
There are several implementations of differential phase-measuring methods. The underlying principle is mixing the backscattered signal with itself at a relative delay and detecting the phase of the resulting cross-correlation product. The two signals can be mixed either in the optical domain (direct detection techniques) or the electrical domain (coherent detection techniques). They can have the same frequency (interferometric recovery technique and homodyne detection) or slightly different frequencies (dual-pulse systems or heterodyne detection).Hartog2017 ; He2021
In all cases, the resulting interferometric field is detected with coherent optical reflectometry methods, Eq. (41), by mixing the backscattered light $\widetilde{E}_{b}(\tau)$ with a reference signal, which is the corresponding backscattered light at a relative delay of $\Delta\tau$, $E_{\text{ref}}=\widetilde{E}_{b}(\tau+\Delta\tau)$. For an undisturbed fiber, this delay corresponds to the constant gauge length $L_{g}$ along the fiber. External mechanical vibrations deform the optical fiber cable and perturb the gauge length by $\Delta L$ resulting in a relative delay of $\Delta\tau=2(L_{g}+\Delta L)/c_{n}$. Hence, the interference product at the photodetector is [Hartog2017, , Eq. (6.11)]
$\widetilde{E}_{b}\left(\tau\right)\widetilde{E}^{*}_{b}\left(\tau+\Delta\tau\right)$ (20)
$=\Big{(}e^{j\phi_{0}}\!\!\!\!\!\int\limits_{z-L_{p}/2}^{z+L_{p}/2}\!\!\!\!\!w\left(z-z_{s}\right)e^{j\omega_{p}\left(\tau-2\frac{z_{s}}{c_{n}}\right)}s(z_{s})\mathrm{d}z_{s}\Big{)}$
$\cdot\Big{(}e^{j\phi_{0}}\!\!\!\!\!\!\!\!\!\!\!\!\!\!\int\limits_{z+L_{g}+\Delta L-L_{p}/2}^{z+L_{g}+\Delta L+L_{p}/2}\!\!\!\!\!\!\!\!\!\!\!\!\!\!w\left(z\!+\!L_{g}+\Delta L\!-\!z_{s}\right)e^{j\omega_{p}\left(\tau+\Delta\tau-2\frac{z_{s}}{c_{n}}\right)}s(z_{s})\mathrm{d}z_{s}\Big{)}^{*},$
or expressed in the spatial domain,
$\widetilde{E}_{b}\left(z\right)\widetilde{E}^{*}_{b}\left(z+L_{g}+\Delta L\right)$ (21)
$=\!e^{j(2k_{p}z+\phi_{0})}\!\!\!\!\!\int\limits_{z-L_{p}/2}^{z+L_{p}/2}\!\!\!\!w(z\!-\!z_{s})e^{-j2k_{p}z_{s}}s(z_{s})\mathrm{d}z_{s}$
$\cdot\!e^{-j(2k_{p}(z+L_{g}+\Delta L)+\phi_{0})}\Big{(}\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\int\limits_{z+L_{g}+\Delta L-L_{p}/2}^{z+L_{g}+\Delta L+L_{p}/2}\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!w(z\!-\!z_{s})e^{-j2k_{p}z_{s}}s(z_{s})\mathrm{d}z_{s}\!\Big{)}^{*},$
with wavenumber spectrum,
$\widetilde{E}_{b}\left(k\right)\widetilde{E}^{*}_{b}(k)$ $=e^{-j2k_{p}(L_{g}+\Delta L)}$ (22)
$\quad\cdot\left[\widetilde{W}(k)\widetilde{S}(k\!-\!k_{p})\!\right]\!\!\left[\widetilde{W}(k)\widetilde{S}(k\!-\!k_{p})\!\right]^{*}$
$=e^{-j2k_{p}(L_{g}+\Delta L)}|\widetilde{W}(k)|^{2}|\widetilde{S}(k-k_{p})|^{2}$
@@
where $\Delta\Phi=2k_{p}\Delta L$ is the differential phase and $\widetilde{C}_{b(L_{g})}(k)=|\widetilde{W}(k)|^{2}|\widetilde{S}(k-k_{p})|^{2}e^{-j2k_{p}L_{g}}$ denotes the cross-spectral density of the backscattered signal at a relative delay determined by the gauge length $L_{g}$.
The gauge length determines the amount of overlap of the backscattered signals in Eq. (20), i.e., $L_{\text{overlap}}=\max\left(L_{p}-L_{g},0\right)$. The mixed signals are correlated for a gauge length shorter than the pulse length, $L_{g}<L_{p}$. Hence, the gauge length should be longer than the pulse length to measure the effect of phase modulation due to external vibrations independently of the pulse that probes the fiber.
Figure 4: Three-dimensional Cartesian coordinate system indicating the direction $\mathbf{e}$ of an impinging wavefield in terms of the angle $\phi$ defined anti-clockwise between the $x$ and the $y$-axis and the angle $\theta$ defined anti-clockwise between the $xy$-plane and the $z$-axis. By convention, the linear optical fiber cable is aligned with the z-axis. The two-dimensional perspective shows that the angle $\theta$ corresponds to the wavefield’s angle of incidence.
Figure 5: Two-dimensional data structure in DAS [adapted from Hartog2017, , Fig. 6.1]. (a) An interrogator uses a laser source to transmit a pulse train $E_{p}(t)$ into the optical fiber, and (b) receives the backscattered signal $E_{b}(t)$ due to inhomogeneities in the dielectric material. (c) Back at the laser’s position, the received time series is converted to slow time $t$ and fast time $\tau$, which corresponds to the two-way travel time of the backscattered wave along the fiber. The time scales are determined by the acquisition sampling frequency $f_{s}$ along the fast-time axis $\tau$, and the pulse repetition frequency $f_{r}$ along the slow-time axis $t$. (d) Measurements along the fast-time correspond to the spatial dimension along the fiber, $z=\tau c_{n}/2$, and are commonly downsampled to the channel spacing $L_{h}$. Phase variations in the backscattered signal measured along distance $z$ from the interoggator are associated with the applied axial strain $\epsilon_{zz}$, whereas phase variations in the backscattered signal measured across time $t$ are associated with the strain rate $\partial\epsilon_{zz}/\partial t$.
@@ II.5 Axial strain estimates
The following analysis shows that the differential phase measurements of the light backscatter (optics) are linearly related to the axial fiber strain (mechanics), which, in turn, is non-linearly related to acoustic quantities in the environment (acoustics). The subscript $a$ is introduced to indicate acoustic quantities and differentiate them from the optical quantities in the previous analysis. Namely, $\omega_{a}$ denotes the angular frequency of the acoustic wave, $k_{a}=\omega_{a}/c_{a}$ denotes the wavenumber, and $c_{a}$ denotes the sound speed.
Interrogating the optical fiber cable with phase-sensitive coherent reflectometry methods provides a distributed measurement of the differential phase $\Delta\Phi$ caused by optical path length variations $\Delta L$; see Eq. (22). Specifically,
$\Delta\Phi$ $=2k_{p}\Delta L=2\frac{2\pi(n+\Delta n)}{\lambda_{p}}L_{g}\frac{\Delta L}{L_{g}}$ (23)
$=\frac{4\pi n\xi L_{g}}{\lambda_{p}}\frac{\Delta L}{L_{g}},$
where $\lambda_{p}$ is the wavelength of the probe pulse and $\xi=1+\Delta n/n$ is the Pockels coefficient that accounts for the strain-optical effect, i.e., strain applied to the fiber both changes its length and modifies its refractive index.Hartog2017 ; williams2024toward In single-mode fiberglass, the Pockels coefficient is $\xi=0.79$.Lindsey2021 Note that $\Delta n$ describes a change in the refractive index due to strain, not the differential change $\mathrm{d}n$ due to material inhomogeneities in Eq. (1).
Considering an optical fiber cable that is coupled with an elastic medium, e.g., deployed on the seafloor, the cable’s deformation $\Delta L$ is attributed to mechanical waves impinging on it [Fig. 2(e)]. Specifically, the particle displacement vector in three-dimensional space $\mathbf{r}=[x,y,z]^{T}$, due to an acoustic plane wave with amplitude $A$ is expressed as,
$\widetilde{\mathbf{u}}(t,\mathbf{r})=(A\mathbf{e})\odot e^{j(\omega_{a}t-k_{a}\mathbf{e}\odot\mathbf{r})},$ (24)
where the direction of the plane wave is defined by the unit vector $\mathbf{e}$ and the $\odot$ operator denotes element-wise multiplication.
The directional unit vector in Cartesian coordinates is defined as $\mathbf{e}=[\cos\phi\cos\theta,\sin\phi\cos\theta,\sin\theta]^{T}$, where $\phi$ is the anti-clockwise angle between the $x$ and $y$-axis and $\theta$ is the anti-clockwise angle between the $xy$-plane and the $z$-axis, as depicted in Fig. 4. By convention, the linear optical fiber cable is aligned with the $z$-axis, such that the angle $\theta$ corresponds to the incident angle of the acoustic plane wave. Alternatively, the direction of the plane wave can be defined in terms of the grazing angle $\theta_{g}=90^{\circ}-\theta$, i.e., the complementary of the incident angle, such that $\mathbf{e}=[\cos\phi\sin\theta_{g},\sin\phi\sin\theta_{g},\cos\theta_{g}]^{T}$.
The particle displacement along the $z$ direction, $\widetilde{u}_{z}(t,z)=A\sin\theta e^{j(\omega_{a}t-k_{a}z\sin\theta)}$, is related to the axial strain as [Ewing1957, , Eq. (1-2)]
@@
$=\widetilde{u}_{z}(t,z)\frac{e^{-jk_{a}\frac{L_{g}}{2}\sin\theta}-e^{jk_{a}\frac{L_{g}}{2}\sin\theta}}{L_{g}}$
$=\widetilde{u}_{z}(t,z)\frac{-j2\sin(k_{a}\frac{L_{g}}{2}\sin\theta)}{L_{g}}$
$=\widetilde{u}_{z}(t,z)\left(-jk_{a}\sin\theta\right)\frac{\sin(k_{a}\frac{L_{g}}{2}\sin\theta)}{k_{a}\frac{L_{g}}{2}\sin\theta}$
$=\!\!-jk_{a}\sin^{2}\!\theta Ae^{j(\omega_{a}t-k_{a}z\sin\!\theta)}\mathrm{sinc}(k_{z}L_{g}/2),$
where $k_{z}=k_{a}\sin\theta$ and the spatial derivative is approximated by the total deformation over a finite gauge length, ${\Delta L}/{L_{g}}$.
Equation (23) indicates that the measured differential phase is linearly related to an estimate of the axial strain over the gauge length $\widehat{\epsilon}_{zz}=\Delta L/L_{g}$ (25) as
$\widehat{\epsilon}_{zz}=\frac{\lambda_{p}}{4\pi n\xi L_{g}}\Delta\Phi.$ (26)
In turn, the axial strain estimate in Eq. (26) is related to the particle displacement of external vibrations such as the acoustic wavefield around the optical-fiber cable through Eq. (25), which reveals two important aspects for interpreting DAS measurements.
First, the dependence of the axial strain amplitude to the square of the sine of the wavefield’s incident angle makes the strain measurement insensitive to broadside incidence, i.e., $\widetilde{\epsilon}_{zz}=0$ for $\theta=0$. This angular dependency is crucial in interpreting DAS data, as it influences the amplitude and detectability of acoustic signals. It also limits the effectiveness of DAS in configurations where ray paths are predominantly orthogonal to the fiber-optic cable. To overcome this limitation, helically wound cables were developed offering broadside sensitivity and enabling the observation of signals from a wider range of angles.kuvshinov2016interaction ; hornman2017field ; al2023experiences Nevertheless, DAS is widely applied on fiber optic cables deployed for telecommunications, which comprise straight optical fibers.
Second, the finite difference approximation of the spatial derivative of the displacement over the gauge length introduces directivity to the measurement defined by the term $\mathrm{sinc}(k_{z}L_{g}/2)$. Hence, the gauge length acts as a spatial window which determines the measurement resolution; see Sec. III for details.
Some DAS systems, instead of estimating the axial strain in Eq. (26) by unwrapping the differential phase measurement $\Delta\Phi$, estimate the strain rate by differentiating the phase measurement between successive pulses [Hartog2017, , Sec. 6.3.2], $\frac{\partial\epsilon_{zz}}{\partial t}$. In such systems, denoting the pulse repetition period as $T$, the axial strain rate estimate is
@@
$\approx(-jk_{a}\sin\theta)\mathrm{sinc}(k_{z}L_{g}/2)\frac{\partial\widetilde{u}_{z}}{\partial t}$
$\approx\omega_{a}k_{a}\sin^{2}\theta Ae^{j(\omega_{a}t-k_{a}z\sin\theta)}\mathrm{sinc}(k_{z}L_{g}/2).$
Concisely, the axial strain is related to the particle displacement component along the fiber, whereas the axial strain rate is related to the corresponding component of particle velocity.